Suppose a 650 kg pallet of boxes is lifted up and stored on a Shelf 2.5 m above-the-ground how much work was done by a forklift to lift the pallet and story on the Shelf
1
Expert's answer
2019-01-18T11:58:01-0500

We can find how much work was done by the forklift from the formula:

"W = mgh = 650 kg \cdot 9.8 \frac{m}{s^2} \cdot 2.5 m = 15925 J."

Answer:

"W = 15925 J"

.